OpenStudy (anonymous):

2 cosxsinx=sinx on [0, 2pi] please find all solutions

OpenStudy (campbell_st):

lots of ways do do this problem here is 1 2cos(x)(sin(x) - sin(x) = 0 take out the common factor sin(x)(2cos(x) - 1) = 0 to make the equation true sin(x) = 0 and cos(x) = 1/2 (1st and 4th quadrants) you just need to find the angles...

OpenStudy (anonymous):

Thanks You! I was wondering if I was doing it right method wise. Thank you again :)

OpenStudy (campbell_st):

if you divide both sides by sin(x) you will miss the solution sin(x) = 0 and only have cos(x) = 1/2

OpenStudy (anonymous):

its 0,pi, pi/3, 5pi/3 right?

OpenStudy (campbell_st):

thats it. well done